Some bank robbers leave town, speeding at 80mph. 15min later, the polive give chase, traveling at 90mph. How long (hrs) will it take the police to overtke the robbers? Can you explain how you get the answer?
Answer by AnlytcPhil(1807)

Some bank robbers leave town, speeding at 80mph. 15min later, the polive give chase, traveling at 90mph. How long (hrs) will it take the police to overtke the robbers? Can you explain how you get the answer?
'
15 minutes after robbers leave town they have traveled 1/4 hour at 80 mph, and thus have traveled 20 miles, (using D = RT = (80 mi/hr)(1/4 hr) = 20 miles.)
'
So the robbers have a 20 mile lead on the cops when the cops set out.
'
Since the cops are going 90 mph and the robbers are going 80 mph, the cops are gaining on the robbers by 10 mph. Therefore the cops will gain the whole 20 mile lead the robbers have on them in 2 hours, (using T = D/R = 20/10 = 2).
'
Edwin
AnlytcPhil@aol.com